# Stable-MoE: Lyapunov-based Token Routing for Distributed Mixture-of-Experts Training over Edge Networks

### v1

@@ I Introduction
Mixture of experts (MoE) models have recently gained significant attention as an effective architecture for large-scale neural networks, especially in distributed training systems where experts are trained or deployed on edge devices [Zhu2025Trustworthy, Zhang2025Communication, xue2025wdmoe]. The main advantage of MoE lies in its sparse activation mechanism, which selectively activates only a subset of experts for each token input. This design significantly reduces computational load and enhances the scalability of the model [fedus2022switch]. A central module of MoE is token routing, which determines the assignment of each token input to the most suitable experts. It is well established that an efficient routing strategy is crucial for optimizing the utilization of computational resources in MoE while simultaneously satisfying the demand of high-quality learning performance [shazeer2017outrageously].
@@
In this context, various token routing strategies have been developed in centralized MoE architectures. For example, a pioneering work of MoE in [shazeer2017outrageously] employs a simple top-$K$ gating mechanism that routes each token to $K$ experts with the highest gating scores based on token feature. Later, Switch Transformer of [fedus2022switch] simplifies this approach by routing each token to the best expert only, while GLaM of [du2022glam] uses the top-$2$ routing strategy combined with auxiliary load balancing losses. GShard of [lepikhin2021gshard] also adopts the top-$2$ routing strategy but enforces expert capacity constraints by discarding tokens that exceed the per-expert token quota. Specifically, this quota refers to the maximum number of tokens that each expert is allowed to process per batch. Moving forward, BASE Layers of [lewis2021base] reformulates routing as a balanced assignment problem to ensure that tokens are evenly distributed across experts without being discarded. Recent works such as Expert Choice Routing of [zhou2022mixture] address the issues of expert utilization by inverting the traditional routing paradigm: instead of tokens being routed to experts based on their compatibility, experts actively select tokens that align with their specialized capabilities.
@@
As the deployment of MoE becomes promising in edge networks for distributed training, traditional token routing strategies face new challenges specific to resource-heterogeneous environments [zhu2025moe, Zhu2025Randomized]. Specifically, in distributed MoE training, routing decisions are required to account for diverse resource constraints and random channel. For instance, the work of [xue2025wdmoe] proposes WDMoE to mitigate delays in wireless networks that considers end-to-end latency into token routing. Beyond the latency-aware routing, recent works such as SMILE of [he2022smile] explore bi-level routing to adapt to heterogeneous network bandwidths, where tokens are first routed to a computing node and then to a specific GPU within that node. Later, the work of [song2025mixture] proposes channel-aware gating functions that incorporate wireless communication quality into token routing for edge computing scenarios. Despite the aforementioned routing strategies, a critical end-to-end throughput bottleneck manifests in practical deployments: under dynamic token arrival, token backlogs may occur at the edge devices with heterogeneous computational capacities, which in turn results in pronounced workload backlog and performance degradation. To date, how to address this token backlog issue remains an open and non-trivial research problem.
@@ II-A System Model
Fig. 1 illustrates a distributed MoE architecture over an edge network consisting of a router and $J$ heterogeneous edge servers, wherein the gating network is executed in the router and different experts are distributed over different edge servers [xue2025wdmoe]. The distributed MoE training over edge networks proceeds as follows:
@@
where the floor function $\lfloor\cdot\rfloor$ rounds a number down to the nearest integer, $\tau_{j}^{\mathrm{com}}=c_{i}/f_{j}(t)$ is the computation time required by edge server $j$ to process a single token, $d_{j}^{\mathrm{rou}}(t)=\sum_{i\in\mathcal{S}(t)}x_{i,j}(t)$ is the total number of tokens routed to the expert $j$, $c_{i}$ is the number of CPU cycles for computing every single token, and $f_{j}(t)$ is the operating CPU frequency of edge server $j$. Note that the duration of each time slot should be significantly longer than the computation time of a single token, i.e., $0\leq\tau_{j}^{\mathrm{com}}(t)\leq\tau$. The CPU frequency is constrained by the edge server’s maximum computational capability $0\leq f_{j}(t)\leq f^{\mathrm{max}}.$ Let $Q_{j}(t)$ be the backlog of the token queue at edge server $j$ over the time slot $t$. The token queuing dynamics at edge server $j$ is given by [neely2010stochastic]
@@
At each time slot, each edge server incurs energy consumption due to executing the routed computational tokens. Specifically, the computation energy consumed by edge server $j$ at the time slot $t$ is given by [mao2016dynamic, deng2020wireless]
@@
where $\xi_{j}$ is the effective switched capacitance determined by the edge server’s chip architecture [burd1996processor]. In practice, the energy budget of each edge server is limited, i.e., $0\leq E_{j}^{\mathrm{com}}(t)\leq E_{j}^{\mathrm{max}}$. From (1) and (3), the computational capabilities (i.e., the maximum number of tokens computed per time slot) of different edge servers are diverse. To manage and constrain the energy usage of edge servers, we introduce an energy queue that reflects the long-term evolution of energy consumption. The energy queuing dynamics at edge server $j$ is given by
@@ II-B Problem Formulation
Under traditional top-$K$ routing strategy such as [shazeer2017outrageously], the gating network scores and activates the experts solely based on token features, but neglects the impact of resource constraints and token backlogs of edge servers. This inevitably results in inefficient resource utilization and thereby degrades system throughput, where some servers are overloaded with excessive tokens while the others remain underutilized.
@@
$\mathbf{P}:$ $\max_{\bm{X}(t)}{\gamma(\overline{\bm{d}})+\mu\overline{G}}$ (5)
@@
$\text{C}5:\lim_{t\to\infty}\frac{\mathbb{E}\{|Q_{j}(t)|\}}{t}=0,\forall j\in\mathcal{J},t\in\mathcal{T},$
$\text{C}6:\lim_{t\to\infty}\frac{\mathbb{E}\{|Z_{j}(t)|\}}{t}=0,\forall j\in\mathcal{J},t\in\mathcal{T},$
@@
where $\gamma(\overline{\bm{d}})=\sum_{j\in\mathcal{J}}\mathrm{log}(1+\overline{d_{j}^{\mathrm{com}}})$ is the system utility that positively correlates with the system throughput [neely2010stochastic]. Note that $\gamma(\overline{\bm{d}})$ is a concave and increasing function that balances throughput and fairness by discouraging excessive load on individual edge servers. $\overline{d_{j}^{\mathrm{com}}}=\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[d_{j}^{\mathrm{com}}(t)]$ is the time average of completed tokens at edge server $j$, and $\overline{\bm{d}}=[\overline{d_{1}^{\mathrm{com}}},\ldots,\overline{d_{J}^{\mathrm{com}}}]$. We optimize $G(t)=\sum_{i\in\mathcal{S}(t)}\sum_{j\in\mathcal{J}}g_{i,j}(t)x_{i,j}(t)$ to maximize the consistency between traditional and optimized routing strategies, where $g_{i,j}(t)\in[0,1]$ is the gating scores of traditional MoE. Additionally, $\overline{G}=\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[G(t)]$ is the time average of $G(t)$, and $\mu\geq 0$ serves as a weighting factor that controls the trade-off between system throughput and gating consistency.
@@
Notably, problem P is a stochastic optimization problem that aims to jointly optimize long-term system throughput and routing consistency under stochastic token arrivals and energy constraints. Specifically, the objective function consists of a logarithmic throughput utility term and a soft regularization term for gating consistency. Constraints ($\text{C}1$)-($\text{C}4$) have been defined in Section II-A. Moreover, $\text{C}5$ and $\text{C}6$ ensure long-term stability of the token and energy queues, respectively [neely2010stochastic].
III Proposed Stable-MoE
@@
It is rather challenging to solve P offline due to coupled decisions, long-term objectives, and system uncertainty. To address this, Lyapunov optimization [neely2010stochastic] offers a principled framework for making sequential decisions in dynamic systems with time-varying states and unknown future information. In this paper, we use the Lyapunov method to transform the intractable long-term optimization problem into tractable subproblems per time slot.
@@
For the queuing dynamics $Q_{j}(t)$ and $Z_{j}(t)$, we define the Lyapunov function as $L(t)=\frac{1}{2}\{\sum_{j\in\mathcal{J}}(Q_{j}(t)^{2}+Z_{j}(t)^{2})\}$. Let $\bm{\Theta}(t)=\{Q_{j}(t),Z_{j}(t),\forall j\in\mathcal{J}\}$ denote the system state at time slot $t$. We define the conditional Lyapunov drift for time slot $t$ as $\Delta(t)=\mathbb{E}[L(t+1)-L(t)|\bm{\Theta}(t)]$, which depends on the scheduling policy in response to dynamically arriving tokens and the current system state $\bm{\Theta}(t)$. Minimizing $\Delta(t)$ helps stabilize the virtual queues $\bm{\Theta}(t)$, thereby promoting satisfaction of the long-term stability constraints $\text{C}5$ and $\text{C}6$ [neely2010stochastic]. However, minimizing $\Delta(t)$ alone may result in low system throughput. Instead, we minimize the Lyapunov drift-plus-penalty function $\Delta_{V}(t)=\Delta(t)-V\mathbb{E}[\gamma(\overline{\bm{d}})+\mu\overline{G}|\bm{\Theta}(t)]$, where $V>0$ is a tunable parameter that controls the trade-off between optimization objective and queue stability.
@@
$\frac{1}{2}\sum_{j\in\mathcal{J}}[Q_{j}(t+1)^{2}-Q_{j}(t)^{2}]$ (9)
$\leq\sum_{j\in\mathcal{J}}Q_{j}(t)[d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t)]+\frac{1}{2}\sum_{j\in\mathcal{J}}[d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t)]^{2}.$
Using the similar manipulations for the energy queue in (4), we have
$\frac{1}{2}\sum_{j\in\mathcal{J}}[Z_{j}(t+1)^{2}-Z_{j}(t)^{2}]$ (10)
$\leq\sum_{j\in\mathcal{J}}Z_{j}(t)[E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}}]+\frac{1}{2}\sum_{j\in\mathcal{J}}[E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}}]^{2}.$
Summing up (9) and (10), taking the conditional expectation, we have $\Delta(t)\leq\sum_{j\in\mathcal{J}}Q_{j}(t)\mathbb{E}[d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t)|\bm{\Theta}(t)]+\sum_{j\in\mathcal{J}}Z_{j}(t)\mathbb{E}[E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}}|\bm{\Theta}(t)]+\sum_{j\in\mathcal{J}}\frac{1}{2}\mathbb{E}[(d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t))^{2}+(E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}})^{2}|\bm{\Theta}(t)].$ Then, subtracting $V\mathbb{E}[\sum_{j=1}^{J}\mathrm{log}(1+d_{j}^{\mathrm{com}}(t))+\mu\sum_{i=1}^{|\mathcal{S}(t)|}\sum_{j=1}^{J}g_{i,j}(t)x_{i,j}(t)|\bm{\Theta}(t)]$ from both sides yields
@@
Note that $d_{j}^{\mathrm{rou}}(t),d_{j}^{\mathrm{com}}(t),$ and $E_{j}^{\mathrm{com}}(t)$ are non-negative. Then, by $(a-b)^{2}\leq a^{2}+b^{2}$ for non-negative $a$ and $b$, we have $\frac{1}{2}\mathbb{E}[(d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t))^{2}|\bm{\Theta}(t)]\leq\frac{1}{2}\mathbb{E}[d_{j}^{\mathrm{rou}}(t)^{2}]+\frac{1}{2}\mathbb{E}[d_{j}^{\mathrm{com}}(t)^{2}].$ Since $0\leq d_{j}^{\mathrm{rou}}(t)\leq|\mathcal{S}(t)|$, it follows that $\mathbb{E}[d_{j}^{\mathrm{rou}}(t)^{2}]\leq\mathbb{E}[\mathcal{S}(t)^{2}]=\lambda+\lambda^{2}$. Moreover, let $D_{j}^{\mathrm{max}}=\lfloor\frac{\tau f_{j}^{\mathrm{max}}}{c_{i}}\rfloor$ denote the maximum processing capacity of edge server $j$, we obtain $\mathbb{E}[d_{j}^{\mathrm{com}}(t)^{2}]\leq(D_{j}^{\mathrm{max}})^{2}$. Similarly, $\mathbb{E}[(E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}})^{2}|\bm{\Theta}(t)]\leq(E_{j}^{\mathrm{max}})^{2}+(E_{j}^{\mathrm{avg}})^{2}$. Thus, we arrive at (6). ∎
Note that the upper bound of $\Delta_{V}(t)$ involves conditional expectations. To minimize the upper bound of $\Delta_{V}(t)$, we employ opportunistic expectation minimization of [neely2010stochastic] to generate the optimal policy. Specifically, in each time slot, we generate an implementable scheduling policy to minimize
@@
$\mathbf{P1}:\max_{\mathbf{X}(t)}$ $V[\sum_{j\in\mathcal{J}}\mathrm{log}(1+d_{j}^{\mathrm{com}}(t))+\mu\sum_{i\in\mathcal{S}(t)}\sum_{j\in\mathcal{J}}g_{i,j}(t)$
$x_{i,j}(t)]-\{\sum_{j\in\mathcal{J}}Q_{j}(t)[d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t)]+\sum_{j\in\mathcal{J}}$
$Z_{j}(t)[E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}}]\}$
s.t. $\text{C}1\sim\text{C}4.$ (13)
As discussed above, the per-slot optimization problem P1 involves both integer decision variables and a nonlinear objective function. Due to the complex coupling among variables, this problem resists analytical solutions and efficient decomposition. Therefore, we employ a solver-based approach based on mixed-integer programming techniques (e.g., branch-and-bound) to optimize the token routing matrix and computation frequency allocation at each time slot. The pseudocode of Stable-MoE is summarized in Algorithm 1.

### v2

@@ I Introduction
Mixture of experts (MoE) models have recently gained significant attention as an effective architecture for large-scale neural networks, especially in distributed training systems where experts are trained or deployed on edge devices [19, 17, 13]. The main advantage of MoE lies in its sparse activation mechanism, which selectively activates only a subset of experts for each token input. This design significantly reduces computational load and enhances the scalability of the model [5]. A central module of MoE is token routing, which determines the assignment of each token input to the most suitable experts. It is well established that an efficient routing strategy is crucial for optimizing the utilization of computational resources in MoE while simultaneously satisfying the demand of high-quality learning performance [14].
@@
In this context, various token routing strategies have been developed in centralized MoE architectures. For example, a pioneering work of MoE in [14] employs a simple top-$K$ gating mechanism that routes each token to $K$ experts with the highest gating scores based on token feature. Later, Switch Transformer of [5] simplifies this approach by routing each token to the best expert only, while GLaM of [4] uses the top-$2$ routing strategy combined with auxiliary load balancing losses. GShard of [9] also adopts the top-$2$ routing strategy but enforces expert capacity constraints by discarding tokens that exceed the per-expert token quota. Specifically, this quota refers to the maximum number of tokens that each expert is allowed to process per batch. Recent works such as Expert Choice Routing of [18] address the issues of expert utilization by inverting the traditional routing paradigm: instead of tokens being routed to experts based on their compatibility, experts actively select tokens that align with specialized capabilities.
@@
As MoE becomes promising in distributed edge intelligence, its implementation faces new challenges in terms of trustworthiness [20], security [2, 7], and privacy [16]. Beyond these, routing decisions of MoE over wireless networks are required to account for diverse resource constraints and random channel conditions. For instance, the work of [17] proposes WDMoE to mitigate delays in wireless networks that considers end-to-end latency into token routing. Beyond the latency-aware routing, recent works such as SMILE of [6] explore bi-level routing to adapt to heterogeneous network bandwidths, where tokens are first routed to a computing node and then to a specific GPU within that node. Later, the work of [15] proposes channel-aware gating functions that incorporate wireless communication quality into token routing for edge computing scenarios. Despite the aforementioned routing strategies, a critical end-to-end throughput bottleneck manifests in practical deployments: under dynamic token arrival, token backlogs may occur at the edge devices with heterogeneous computational capacities, which in turn results in pronounced workload backlog and performance degradation. To date, how to address this token backlog issue remains an open and non-trivial research problem.
@@ II-A System Model
Fig. 1 illustrates a distributed MoE architecture over an edge network consisting of a router and $J$ heterogeneous edge servers, wherein the gating network is executed in the router and different experts are distributed over different edge servers [17]. The distributed MoE training over edge networks proceeds as follows:
@@
where the floor function $\lfloor\cdot\rfloor$ rounds a number down to the nearest integer, $\tau_{j}^{\mathrm{com}}=c_{i}/f_{j}(t)$ is the computation time required by edge server $j$ to process a single token, $d_{j}^{\mathrm{rou}}(t)=\sum_{i\in\mathcal{S}(t)}x_{i,j}(t)$ is the total number of tokens routed to the expert $j$, $c_{i}$ is the number of CPU cycles for computing every single token, and $f_{j}(t)$ is the operating CPU frequency of edge server $j$. Note that the duration of each time slot should be significantly longer than the computation time of a single token, i.e., $0\leq\tau_{j}^{\mathrm{com}}(t)\leq\tau$. The CPU frequency is constrained by the edge server’s maximum computational capability $0\leq f_{j}(t)\leq f^{\mathrm{max}}.$ Let $Q_{j}(t)$ be the backlog of the token queue at edge server $j$ over the time slot $t$. The token queuing dynamics at edge server $j$ is given by [11]
@@
At each time slot, each edge server incurs energy consumption due to executing the routed computational tokens. Specifically, the computation energy consumed by edge server $j$ at the time slot $t$ is given by [10, 3]
@@
where $\xi_{j}$ is the effective switched capacitance determined by the edge server’s chip architecture [1]. In practice, the energy budget of each edge server is limited, i.e., $0\leq E_{j}^{\mathrm{com}}(t)\leq E_{j}^{\mathrm{max}}$. From (1) and (3), the computational capabilities (i.e., the maximum number of tokens computed per time slot) of different edge servers are diverse. To manage and constrain the energy usage of edge servers, we introduce an energy queue that reflects the long-term evolution of energy consumption. The energy queuing dynamics at edge server $j$ is given by
@@ II-B Problem Formulation
Under the traditional top-$K$ routing strategy such as [14], the gating network scores and activates the experts solely based on token features, but neglects the impact of resource constraints and token backlogs of edge servers. This inevitably results in inefficient resource utilization and thereby degrades system throughput, where some servers are overloaded with excessive tokens while the others remain underutilized.
@@
$\mathbf{P}:$ $\max_{\bm{X}(t)}\kern 5.0pt{\gamma(\overline{\bm{d}})+\mu\overline{G}}$ (5)
@@
$\text{C}5:\lim_{t\to\infty}\mathbb{E}\{|Q_{j}(t)|\}/{t}=0,\forall j\in\mathcal{J},t\in\mathcal{T},$
$\text{C}6:\lim_{t\to\infty}\mathbb{E}\{|Z_{j}(t)|\}/{t}=0,\forall j\in\mathcal{J},t\in\mathcal{T},$
@@
where $\gamma(\overline{\bm{d}})=\sum_{j\in\mathcal{J}}\mathrm{log}(1+\overline{d_{j}^{\mathrm{com}}})$ is the system utility that positively correlates with the system throughput [11]. Note that $\gamma(\overline{\bm{d}})$ is a concave and increasing function that balances throughput and fairness by discouraging excessive load on individual edge servers. $\overline{d_{j}^{\mathrm{com}}}=\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[d_{j}^{\mathrm{com}}(t)]$ is the time average of completed tokens at edge server $j$, and $\overline{\bm{d}}=[\overline{d_{1}^{\mathrm{com}}},\ldots,\overline{d_{J}^{\mathrm{com}}}]$. We optimize $G(t)=\sum_{i\in\mathcal{S}(t)}\sum_{j\in\mathcal{J}}g_{i,j}(t)x_{i,j}(t)$ to maximize the consistency between traditional and optimized routing strategies, where $g_{i,j}(t)\in[0,1]$ is the gating scores of traditional MoE. Specifically, a larger $G(t)$ ensures that tokens are preferentially routed to experts with higher gating scores, thereby preserving model quality while dynamically adapting to the underlying resource constraints. Additionally, $\overline{G}=\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[G(t)]$ is the time average of $G(t)$, and $\mu\geq 0$ serves as a weighting factor that controls the trade-off between system throughput and gating consistency.
@@
Notably, problem P is a stochastic optimization problem that aims to jointly optimize long-term system throughput and routing consistency under stochastic token arrivals and energy constraints. Constraints ($\text{C}1$)-($\text{C}4$) have been defined in Section II-A. Moreover, $\text{C}5$ and $\text{C}6$ ensure long-term stability of the token and energy queues, respectively [11].
III Stable-MoE
@@
It is rather challenging to solve P offline due to coupled decisions, long-term objectives, and system uncertainty. To address this, Lyapunov optimization [11] offers a principled framework for making sequential decisions in dynamic systems with time-varying states and unknown future information. In this paper, we use the Lyapunov method to transform the intractable long-term optimization problem into tractable subproblems per time slot.
@@
For the queuing dynamics $Q_{j}(t)$ and $Z_{j}(t)$, we define the Lyapunov function as $L(t)=\frac{1}{2}\{\sum_{j\in\mathcal{J}}(Q_{j}(t)^{2}+Z_{j}(t)^{2})\}$. Let $\bm{\Theta}(t)=\{Q_{j}(t),Z_{j}(t),\forall j\in\mathcal{J}\}$ denote the system state at time slot $t$. We define the conditional Lyapunov drift for time slot $t$ as $\Delta(t)=\mathbb{E}[L(t+1)-L(t)|\bm{\Theta}(t)]$, which depends on the scheduling policy in response to dynamically arriving tokens and the current system state $\bm{\Theta}(t)$. Minimizing $\Delta(t)$ helps stabilize the virtual queues $\bm{\Theta}(t)$, thereby promoting satisfaction of the long-term stability constraints $\text{C}5$ and $\text{C}6$ [11]. However, minimizing $\Delta(t)$ alone may result in low system throughput. Instead, we minimize the Lyapunov drift-plus-penalty function $\Delta_{V}(t)=\Delta(t)-V\mathbb{E}[\gamma(\overline{\bm{d}})+\mu\overline{G}|\bm{\Theta}(t)]$, where $V>0$ is a tunable parameter that controls the trade-off between optimization objective and queue stability.
@@
$\frac{1}{2}\sum_{j\in\mathcal{J}}[Q_{j}(t+1)^{2}-Q_{j}(t)^{2}]\leq\sum_{j\in\mathcal{J}}Q_{j}(t)[d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t)]$
$+\frac{1}{2}\sum_{j\in\mathcal{J}}[d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t)]^{2}.$ (9)
Similarly, for the energy queue in (4), we have
$\frac{1}{2}\sum_{j\in\mathcal{J}}[Z_{j}(t+1)^{2}-Z_{j}(t)^{2}]\leq\sum_{j\in\mathcal{J}}Z_{j}(t)[E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}}]$
$+\frac{1}{2}\sum_{j\in\mathcal{J}}[E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}}]^{2}.$ (10)
Summing up (III) and (III), taking the conditional expectation, we have $\Delta(t)\leq\sum_{j\in\mathcal{J}}Q_{j}(t)\mathbb{E}[d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t)|\bm{\Theta}(t)]+\sum_{j\in\mathcal{J}}Z_{j}(t)\mathbb{E}[E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}}|\bm{\Theta}(t)]+\sum_{j\in\mathcal{J}}\frac{1}{2}\mathbb{E}[(d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t))^{2}+(E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}})^{2}|\bm{\Theta}(t)].$ Then, subtracting $V\mathbb{E}[\sum_{j=1}^{J}\mathrm{log}(1+d_{j}^{\mathrm{com}}(t))+\mu\sum_{i=1}^{|\mathcal{S}(t)|}\sum_{j=1}^{J}g_{i,j}(t)x_{i,j}(t)|\bm{\Theta}(t)]$ from both sides yields
@@
Note that $d_{j}^{\mathrm{rou}}(t),d_{j}^{\mathrm{com}}(t),$ and $E_{j}^{\mathrm{com}}(t)$ are non-negative. Thus, we have $\frac{1}{2}\mathbb{E}[(d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t))^{2}|\bm{\Theta}(t)]\leq\frac{1}{2}\mathbb{E}[d_{j}^{\mathrm{rou}}(t)^{2}|\bm{\Theta}(t)]+\frac{1}{2}\mathbb{E}[d_{j}^{\mathrm{com}}(t)^{2}|\bm{\Theta}(t)].$ Since $0\leq d_{j}^{\mathrm{rou}}(t)\leq|\mathcal{S}(t)|$, it follows that $\mathbb{E}[d_{j}^{\mathrm{rou}}(t)^{2}|\bm{\Theta}(t)]\leq\mathbb{E}[\mathcal{S}(t)^{2}]=\lambda+\lambda^{2}$. Moreover, let $D_{j}^{\mathrm{max}}=\lfloor\frac{\tau f_{j}^{\mathrm{max}}}{c_{i}}\rfloor$ be the maximum processing capacity of edge server $j$, we obtain $\mathbb{E}[d_{j}^{\mathrm{com}}(t)^{2}|\bm{\Theta}(t)]\leq(D_{j}^{\mathrm{max}})^{2}$. Similarly, $\mathbb{E}[(E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}})^{2}|\bm{\Theta}(t)]\leq(E_{j}^{\mathrm{max}})^{2}+(E_{j}^{\mathrm{avg}})^{2}$. Thus, we arrive at (6). ∎
Note that the upper bound of $\Delta_{V}(t)$ involves conditional expectations. To minimize the upper bound of $\Delta_{V}(t)$, we employ opportunistic expectation minimization of [11] to generate the optimal policy. Specifically, in each time slot, we generate an implementable scheduling policy to minimize
@@
$\mathbf{P1}:\max_{\mathbf{X}(t)}$ $\kern 5.0ptV\Gamma_{\text{u}}(\mathbf{X}(t))-\Gamma_{\text{p}}(\mathbf{X}(t))$
s.t. $\text{C}1\sim\text{C}4,$ (13)
where
$\Gamma_{\text{u}}(\mathbf{X}(t))=\sum_{j\in\mathcal{J}}\mathrm{log}(1+d_{j}^{\mathrm{com}}(t))$
$+\mu\sum_{i\in\mathcal{S}(t)}\sum_{j\in\mathcal{J}}g_{i,j}(t)x_{i,j}(t),$ (14)
$\Gamma_{\text{p}}(\mathbf{X}(t))=\sum_{j\in\mathcal{J}}Q_{j}(t)[d_{j}^{\mathrm{rou}}(t)-d_{j}^{\mathrm{com}}(t)]$
$+\sum_{j\in\mathcal{J}}Z_{j}(t)[E_{j}^{\mathrm{com}}(t)-E_{j}^{\mathrm{avg}}].$ (15)
As discussed above, the per-slot optimization problem P1 involves both integer decision variables and a nonlinear objective function. Due to the complex coupling among variables, we employ a tunable solver-based approach based on mixed-integer programming methods (e.g., branch-and-bound) to optimize the token routing and computation frequency allocation at each time slot. The pseudocode of Stable-MoE is summarized in Algorithm 1.